OpenStudy (anonymous):

Which value(s) of x should be excluded for the polynomial? give me one minute and i will put up the equation A. 3 B. 3 and -3 C. 4 D. 4 and -4

OpenStudy (anonymous):

x^2-16/x^2-6x+9

jimthompson5910 (jim_thompson5910):

You cannot divide by zero So the denominator x^2-6x+9 can't be zero If it were, then x^2-6x+9 = 0 Solve this for x x^2-6x+9 = 0 (x-3)^2 = 0 x-3 = 0 x = 3 So when x = 3, the denominator is zero So we must exclude 3 from the domain So the answer is choice A

OpenStudy (anonymous):

If you plug in 3, you get 0 in the denominator. 3 is an extraneous solution.